OpenStudy (anonymous):

How can you tell if an equation's line, if done on a graph, is parallel/perpendicular/same line..?

OpenStudy (campbell_st):

the slope intercept form of a line is y = mx + b m = slope and b = y-intercept when 2 or more lines have the same slope they are parallel y = 2x y = 2x - 10 m = 2 for both lines perpendicular the slopes multiply to give -1 y = 3x y = -1/3 x + 11 -1/3 * 3 = -1 so perpendicular

OpenStudy (campbell_st):

same line, have the same slope and same intercept y = 4x + 6 and 3y = 12x + 18 divide the 2nd line by 3 and you get the 1st line